Four solutions of an acid dissolved in water are sketched below, as if under a microscope so powerful individual atoms could be seen. The same volume of solution is shown in each sketch. Rank the solutions by the strength of the dissolved acid. That is, select 1 under the solution of the strongest acid, 2 under the solution of the next strongest acid, and so on.

Note: Note: =H,O Solution 1 Solution 2 (Choose one) (Choose one) Solution 3 Solution 4 (Choose one) (Choose one)
Answer #1

Strongest acid rank is determined by the no of H3O+ molecule and total no of acid molecule (whether it is dissociated or undissociated) present in the sketch by using following formula.

Strenth of Acid = (No of H3 Omolecule formed - total no of acid molecule present in the solution) x 100

1) Solution 4 = (4/4)*100= 100%

2) Solution 1 = (5/6)*100= 83%

3) Solution 2 = (3/4)*100= 75%

4) Solution 3 = (3/7)*100= 42%